solve nonlinear equation by using substraction and addition method.
1) {x+2y=0 (x-1)^2 + (y-1)^2=5

Answer by ikleyn(52777)

.
Please accept these my notices:

1.   From https://english.stackexchange.com/questions/3640/is-substract-versus-subtract-a-proper-word

    "Subtract" is the word. Though the obsolete word "substract" did exist, any occurrence you see these days 
     is most likely just a common mistake, formed by analogy either with "abstract" or with other languages 
     whose corresponding words do have two 's's. Many recent dictionaries do not list "substract".


2.   Even if to treat your  "substraction"  as  subtraction,  it DOES NOT work  for this system.


3.   Instead,  the  SUBSTITUTION  method works:

      From the linear equation express  x = -2y  and substitute it into the non-linear equation.  You will get

      (-2y-1)^ + (y-1)^2 = 5.

      It is the quadratic equation.  Simplify it to the standard form and solve by any method.


================
So, the final conclusion is that the condition in this post is INCORRECT.

The correct condition is THIS:

    solve nonlinear equation by using the SUBSTITUTION method.
    1) {x+2y=0 (x-1)^2 + (y-1)^2=5


And I just showed you how to solve it . . .


--------------
See the lesson
    - Solving systems of algebraic equations of degree 2 and degree 1
in this site.

You will find there many similar solved systems.